OpenStudy (anonymous):

Confusing... To create a breakfast beverage, pineapple juice in two concentrations, 16% and 30%, must be combined into a solution that will be mixed with another type of juice to produce the beverage. If 50ml gallon of 16% juice is used, how many gallons of the 30% juice must be used to obtain a 20% pineapple juice solution?

OpenStudy (anonymous):

.20(x+)=.16(50)+0(x) I think?

hero (hero):

Total gallons of pineapple juice 50 + x = y Percent gallons of pineapple juice in each concentration .16(50) + .30x = .20y

hero (hero):

Solve the following system of equations: 50 + x = y 8 + .30x = .20y

hero (hero):

50 + x = y 80 + 3x = 2y 100 + 2x = 2y 80 + 3x = 2y 2y = 2y 100 + 2x = 80 + 3x 100 - 80 = 3x - 2x 20 = x